Find the first four terms of the sequence an= n(5n +3)

Answer by richwmiller(17219)

An= n(5n +3)
Plug in 1 for n
an= n(5n +3)
A1=1*(5+3)
and solve
Plug in 2 for n
A2=2*(2*5+3)
and solve
plug in 3 for n
A3=3*(3*5+3)
and solve
plug in 4 for n
A4=4*(4*5+3)
and solve
Good luck.